OpenStudy (anonymous):

In solving the proportion x-3/4 = 1/4 , which of the following would be your first step? a) 4(x- 3) = 1 b) 1(x- 3) = 4(4) c) 4(x- 3) = 4(1) d) 1(x- 3) = 4

OpenStudy (wondermath):

can't you just move the 3/4 to the other side and solve it?

OpenStudy (anonymous):

4(x-3)=4(1), cross multiply

OpenStudy (anonymous):

You could do that, but it's not one of the answer options

OpenStudy (anonymous):

Then the answer is 4 :D ?

OpenStudy (wondermath):

oh u should have added the brackets x-3/4 = 1/4 and (x-3)/4 = 1/4 mean different things

OpenStudy (anonymous):

Ty and i'm sorry :I

OpenStudy (wondermath):

haha np

hero (hero):

Hold on a sec, I h ave it

hero (hero):